A charge, q1, of -2 micro-coulombs is located at the origin, a charge, q2, of -8 micro-coulombs is located at x = 0 cm, y = +5.5 cm, a charge, q3, of +3 micro-coulombs is located at x = 0 cm, y = -5.2 cm, and a charge, q4, of +12 micro-coulombs is located at x = 2.2 cm, y = 0 cm. What is the angle of the total electric force on q2 (charge on the +y axis) in degrees measured counter-clockwise from the +x axis?
Answer #1

ty q2 5.92 cm 5.5cm 68.2 qi q4 -X 2.2 cm +X 5.2 cm -y

Using a trigonometric identity, we have

heta = tan-1 [(5.5 cm) / (2.2 cm)]

heta = tan-1 (2.5)

heta = 68.2 degree

The force on charge q2 due to q4 which will be given as -

F24 = ke |q2| |q4| / r242

F24 = [(9 x 109 Nm2/C2) (8 x 10-6 C) (12 x 10-6 C)] / (0.0592 m)2

F24 = 246.5 N

The x and y components of F24 which are given below as -

F24,x = (246.5 N) cos 68.20

F24,x = 91.5 N

And

F24,y = (246.5 N) sin 68.20

F24,y = 228.8 N

The force on charge q2 due to q1 which will be given as -

F21 = ke |q2| |q1| / r212

F21 = [(9 x 109 Nm2/C2) (8 x 10-6 C) (2 x 10-6 C)] / (0.055 m)2

F21 = 47.6 N

The x and y components of F21 which are given below as -

F21,x = (47.6 N) cos 900

F21,x = 0 N

And

F21,y = (47.6 N) sin 900

F21,y = 47.6 N

The net electric force due to other three charges on q1 which will be given as -

F2 = (F24,x + F21,x) + (F24,y + F21,y)

F2 = [(91.5 N) + (0 N)] + [(228.8 N) + (47.6 N)]

F2 = (91.5 N) + (276.4 N)

Therefore, magnitude of the net electric force due to other three charges on q2 which will be given by -

| F2 | = sqrt{}(91.5 N)2 + (276.4 N)2

| F2 | = sqrt{}84769.21 N2

| F2 | = 291.1 N

The angle of total electric force on charge q2 which will be given as -

From a trigonometry identity, we get

phi = tan-1 [(276.4 N) / (91.5 N)]

phi = tan-1 (3.0207)

phi = 71.6 degree

{ counter-clockwise from the +x axis }